OpenStudy (anonymous):

Choose the procedure you would use to solve the equation 5x + 6 = 41 for x. Subtract six from both sides, and then multiply by five on both sides. Divide both sides by five, and then add six to both sides. Subtract six from both sides, and then divide by five on both sides. Divide both sides by five, and then subtract five from both sides.

OpenStudy (anonymous):

subtract 6 from both sides . 5x+6=41 5x+6-6=41-6 5x=35 what do you think the next step should be ?

OpenStudy (anonymous):

@Deldrickg waiting for your reply !!!

OpenStudy (anonymous):

divide by 5

OpenStudy (anonymous):

exactly !!!!!!! good.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you got the procedure !!!!

OpenStudy (anonymous):

yes